# Seeing through Light and Darkness: Sensor-Physics Grounded Deblurring HDR NeRF from Single-Exposure Images and Events

### 13 Mar 2026 v3

@@ Abstract
Novel view synthesis from low dynamic range (LDR) blurry images, which are common in the wild, struggles to recover high dynamic range (HDR) and sharp 3D representations in extreme lighting conditions. Although existing methods employ event data to address this issue, they ignore the sensor-physics mismatches between the camera output and physical world radiance, resulting in suboptimal HDR and deblurring results. To cope with this problem, we propose a unified sensor-physics grounded NeRF framework for sharp HDR novel view synthesis from single-exposure blurry LDR images and corresponding events. We employ NeRF to directly represent the actual radiance of the 3D scene in the HDR domain and model raw HDR scene rays hitting the sensor pixels as in the physical world. A pixel-wise RGB mapping field is introduced to align the above rendered pixel values with the sensor-recorded LDR pixel values of the input images. A novel event mapping field is also designed to bridge the physical scene dynamics and actual event sensor output. The two mapping fields are jointly optimized with the NeRF network, leveraging the spatial and temporal dynamic information in events to enhance the sharp HDR 3D representation learning. Experiments on the collected and public datasets demonstrate that our method can achieve state-of-the-art deblurring HDR novel view synthesis results with single-exposure blurry LDR images and corresponding events.
@@ 1 Introduction
Recent novel view synthesis works, such as neural radiance fields (NeRF) [31] and 3D gaussian splatting (3DGS) [23], achieve high-fidelity and high-speed novel view synthesis (NVS) with ideal low dynamic range (LDR) sharp images as input, respectively. However, their performance suffers from two limitations: handling extreme lighting scenes with both highlight and low-light areas and robustness to camera motion blur, which are caused by the inherent information loss in the blurred and compressed LDR images with both underexposed and overexposed regions. Some recent image-based NeRF and 3DGS related works utilize multi-exposure LDR images [6, 4, 27] to compensate for the limited dynamic range and model the camera motion [29, 43, 54] to mitigate the influence of motion blur, while they are still susceptible to the requirement of cumbersome multi-exposure capture and severe motion blur in long-exposure conditions.
@@
Event cameras provide a promising sensing modality by asynchronously recording brightness changes in the log domain, thereby preserving high temporal resolution and dynamic range that are typically lost in conventional frame-based cameras [12]. As shown in the upper half of Fig. 1, event data can effectively record spatial-temporal differential information even in blurred, overexposed, and underexposed areas of an LDR image. Motivated by this advantage, recent approaches have explored using event and RGB image (ERGB) data for the deblurring novel view synthesis task [35, 34, 5, 36, 10], while their performance is limited by a lack of grounding in the sensor-physics mismatch between the recorded RGB image and the physical scene brightness. EvHDR-NeRF [6] first incorporates events into the single-exposure HDR and deblurring NeRF reconstruction task. However, its framework largely follows HDR-NeRF [19] by applying the same tone mapping strategy directly to 3D points and extending it to the event-based setting without considering the event generation character.
@@
We argue that achieving accurate and physically consistent 3D representation learning in the ERGB setting requires more than directly extending existing image-based frameworks. In particular, it is essential to explicitly model how scene radiance is transformed into actual sensor measurements through device-specific processes. For RGB cameras, this involves temporal integration and nonlinear camera response functions; for event cameras, this includes contrast thresholding, latency effects, and photometric quantization. Without accounting for these physical transformations, the rendered outputs of the learned model may not align with the supervision signals derived from input images and events, weakening the learning of scene geometry and appearance as shown in the lower half of Fig. 1. Therefore, a sensor-physics grounded modeling of both RGB and event sensing pipelines is critical for fully exploiting the complementary information in ERGB data and achieving robust HDR and deblurring NVS performance.
@@
In this paper, we present a novel framework that exploits Single-Exposure blurry images and corresponding Events to enable sharp HDR NeRF reconstruction (See-NeRF). Specifically, built upon the physical imaging principles of RGB and event sensors, we adopt NeRF to focus on learning the actual radiance of the scene in the HDR domain and introduce two differentiable mapping fields: a pixel-wise RGB mapping field that models exposure integration and camera response, and an event mapping field that accounts for latency and photometric deviations in event generation. These modules align the predicted radiance with the sensor measurements and are jointly optimized with the NeRF network under the supervision of input images and events. By modeling the radiance-to-measurement process explicitly, See-NeRF can accurately reconstruct physically consistent, sharp HDR 3D representations. Experiments on synthetic and real datasets demonstrate that See-NeRF achieves state-of-the-art HDR and deblurring novel view synthesis results. Our contributions are summarized as follows:
@@
2) We introduce a 2D pixel-wise RGB mapping field and a latency-aware photometrically calibrated event mapping field to bridge the discrepancy between physical radiance values and sensor measurements.
3) We collect both synthetic and real datasets for training and testing. The results show that our method achieves the best performance on both ERGB-based deblurring and HDR novel view synthesis tasks.
@@ 2.1 HDR and deblurring novel view synthesis
Neural radiance fields (NeRF) [31] extensively promote the development of novel view synthesis (NVS) tasks and 3D implicit representation learning. 3D gaussian splatting (3DGS) [23] uses gaussian point clouds to explicitly represent 3D scenes and parallelized rasterization workflow to improve the speed of 3D reconstruction and view rendering. However, NeRF and 3DGS can only learn a sharp LDR scene representation with sharp LDR images. To address the HDR NVS, RawNeRF [30] and RawGS [22] aim to reconstruct an HDR 3D representation from noisy HDR raw images. HDR-NeRF [19] and HDR-GS [4] introduce a 3D points tone mapper to transform the HDR 3D representation into the LDR 3D representation, optimized with multi-exposure LDR images. HDR-HexPlane [46] further extends HDR-NeRF to dynamic scenes. GaussHDR [27] accepts an additional context feature as input to improve HDR NVS. Gaussian in the dark [49] uses a CNN-based feature map to enhance the tone mapper learning. To address the Deblurring NVS from blurry images, Deblur-NeRF [29] proposes a deformable sparse kernel module to model the blur kernels. BAD-NeRF [43] and BAD-GS [54] transform the camera poses into SE(3) space and interpolate the poses to recover the camera motion trajectory. DP-NeRF [24] imposes 3D consistency and refines the color with the relationship between depth and blur. Although these methods have achieved promising performance in HDR and deblurring NVS tasks, they still fundamentally suffer from the inherent information loss of the motion-blurred LDR images.
@@ 2.2 ERGB-based novel view synthesis
Event cameras are bio-inspired sensors detecting brightness changes at individual pixels with high temporal resolution [12]. This scene radiance measurement paradigm has higher dynamic range and temporal resolution compared to traditional frame-based RGB cameras, attracting a lot of research in the fields of event-based image deblurring [33, 21, 26, 40, 47], HDR image reconstruction [42, 15, 44, 32, 45, 14, 8], and high-speed HDR video generation [37, 57, 48, 55]. Some recent works explore event-based NVS [20, 39, 28, 28, 29, 11, 53, 13, 51]. While other works integrate both image and event (ERGB) data into the deblurring NVS task, such as NeRF-based works [35, 34, 5, 36] and 3DGS-based works [10, 18, 50] Addressing the limitations of blurry LDR inputs, EvHDR-NeRF [6] first jointly exploits single-exposure images and events to reconstruct high-fidelity HDR radiance fields. However, these methods do not fully consider the gap between sensor output and physical radiance, leading to limited HDR and deblurring NeRF novel view synthesis performance.
@@ 3 Background
In this section, we leverage the principles of event and LDR image generation to explain why events can enhance HDR deblurring NVS from single-exposure LDR blurry images.
@@ 3.1 From scene radiance to LDR images
In an extreme lighting scene with both highlight and low-light areas, a typical RGB camera sensor records the physical scene radiance through a frame-based integral model, generating a raw image with high dynamic range pixel values. To match human perceptual characteristics and reduce storage costs, the raw image is processed by the image signal processor (ISP), applying non-linear ISO gain, RGB demosaicing, tone mapping, gamma compression, white-balance, and quantization to yield an 8-bit LDR image, which is always modeled by a camera response function (CRF) in practice [2]:
where is the raw scene radiance hitting the sensor at time . and are the start and end time of exposure. is the final color value that we can obtain from the LDR image. This integral model and nonlinear CRF inherently obscure the temporal dynamics of and compress the critical HDR spatial radiance differential information in highlight and low-light areas, ineluctably leading to artifacts like motion blur, overexposure (pixel values saturation at 255), and underexposure (pixel values clipping at 0) in the LDR images, respectively.
@@ 3.2 From scene dynamics to events
Event cameras asynchronously detect the change of the scene radiance hitting the sensor in the log domain and trigger an event whenever the change reaches the threshold . (x,y) represents the pixel coordinates and is the trigger time. The change direction is encoded in . This spike model can preserve temporal radiance dynamics and spatial radiance gradient distribution:
which precisely compensates for the temporal and spatial dynamic loss of blurry compressed LDR images (Sec. 3.1).
@@ 3.3 Events for deblurring and HDR NVS
According to Sec. 3.1, the core problem of recovering the sharp HDR scene radiance from LDR blurry images is formulated as the joint estimation of the latent sharp luminance and the CRF . The temporal radiance dynamics of events in Eq. 2 can be used to estimate the latent sharp along the time dimension from blurry as proved in ERGB-based deblurring works [33, 35]. The spatial dynamics of events in Eq. 2 can provide extra radiance differential information for the low-light and highlight areas and expand the dynamic range of the single-exposure LDR image , as shown in Fig. 2, with which the network in Sec. 4.2 can obtain a wider range of as input and infer a CRF curve that is closer to the ground truth.
@@ 4 Method
In this section, we introduce the See-NeRF framework and theoretically prove how it leverages sensor-physics grounded priors to realize better HDR deblurring and NVS with events and single-exposure blurry LDR images.
@@ 4.1 HDR scene representation
Firstly, we use the NeRF network with parameters to represent the actual radiance and density of a 3D point in an extreme lighting scene rather than the camera-processed LDR radiance as in the original NeRF:
@@
Then we use volume rendering to simulate the process of HDR scene rays hitting the imaging sensor at pixel (x,y) corresponding to the camera pose and obtain the latent sharp HDR radiance at time :
@@
4.2 Physically grounded RGB mapping field
To simulate the integral model of the RGB camera, we discretely sample time points within the exposure time to as in [34], and the generated raw pixel value of the RGB imaging sensor during exposure is:
where is weight parameters calculated by time points [34]. To bridge the gap between the raw sensor value and the LDR image calculated by the CRF in Eq. 1, we used three MLPs to fit this function in the log domain for three color channels separately and obtain the estimated LDR pixels:
@@ 4.2 Physically grounded RGB mapping field
As shown in Fig. 4, unlike HDR-NeRF applying the CRF function directly on the radiance of 3D points in Eq. 3, we employ CRF after volume rendering at the pixel-level for the raw pixel values , aligning with the imaging process in the physical world (Fig. 3), which allows to focus on non-linear color tone mapping learning without being affected by the linear volume rendering, while NeRF focuses on learning 3D density and raw radiance without being affected by . Besides, our blur synthesis is more physically grounded with the HDR pixel values (Eq. 6). Therefore, our proposed RGB mapping field can significantly improve the HDR and deblurring NVS results as demonstrated by the experiment in Sec. 5.4.
4.3 Latency-aware and photometrically calibrated event mapping field
@@
where means taking the integer between the input and zero that is closest to the input. However, the actual event generation may deviate from this ideal model with a significant delay in low-light regions [17], and due to the fixed contrast threshold in event generation, the minimum detectable brightness change is lower-bounded, leading to quantifiable errors in radiometric estimation. To bridge the gap between actual radiance dynamics and actual event generation, we propose an event mapping field comprising three core components: Bayer pattern adaptation, temporal delay compensation, and photometric quantity calibration.
@@
Temporal delay compensation: Since the delay of events is determined by the pixel radiance values of the sensor, we use a compensation function to calculate the coefficient:
Then, we use a first-order low-pass filter to synthesize the event latency with the estimated coefficient :
We omit (x,y) in Eq. 10 because the above operation is the same for all pixels. We set and substitute into the event generation model (Eq. 8) to calculate the estimated number of events:
Photometric quantity calibration: Since event acquisition exhibits uncertainty when brightness changes are below the threshold (e.g., minor brightness variations may not trigger events) and our sampled time points may not precisely align with the timestamp of the last triggered event at this pixel, we pre-estimate an offset from the input events by using , and then subtract this offset for the final estimated number of events to compensate this discrepancy:
@@ 4.4 Loss functions and optimization
We use the LDR image loss and event loss to jointly supervise the optimization of HDR scene radiance representation (Eq. 3) and the two camera mapping fields:
@@
We take and optimize a coarse and a fine network simultaneously as in [34]. During testing, we can obtain the novel view sharp HDR image with Eq. 6 and the sharp LDR image with novel exposure with Eq. 7. We set as a trade-off between performance and training time as in [35]. Detailed implementation details and evaluation on are provided in Sec. C and Sec. D of the supplementary.

### 02 Apr 2026 v4

@@ Abstract
Novel view synthesis from low dynamic range (LDR) blurry images, which are common in the wild, struggles to recover high dynamic range (HDR) and sharp 3D representations in extreme lighting conditions. Although existing methods employ event data to address this issue, they ignore the sensor-physics mismatches between the camera output and physical world radiance, resulting in suboptimal HDR and deblurring results. To cope with this problem, we propose a unified sensor-physics grounded NeRF framework for sharp HDR novel view synthesis from single-exposure blurry LDR images and corresponding events. We employ NeRF to directly represent the actual radiance of the 3D scene in the HDR domain and model raw HDR scene rays hitting the sensor pixels as in the physical world. A 2D pixel-wise RGB CRF model is introduced to align the NeRF rendered pixel values with the sensor-recorded LDR pixel values of the input images. A novel event CRF model is also designed to bridge the gap between physical scene dynamics and event sensor output. The two models are jointly optimized with the NeRF network, leveraging the spatial and temporal dynamic information in events to enhance the sharp HDR 3D representation learning. Experiments on the collected and public datasets demonstrate that our method achieves state-of-the-art HDR and deblurring novel view synthesis results with single-exposure blurry LDR images and corresponding events. Our code and datasets are publicly available at https://github.com/iCVTEAM/See-NeRF.
@@ 1 Introduction
Recent novel view synthesis (NVS) works, Neural Radiance Fields (NeRF) [33] and 3D Gaussian Splatting (3DGS) [23], achieve high-fidelity and high-speed NVS results with ideal low dynamic range (LDR) sharp images as input, respectively. However, their performance suffers from two limitations: handling extreme lighting scenes with both highlight and low-light areas and robustness to camera motion blur, which are caused by the inherent information loss in the blurred and compressed LDR images with both underexposed and overexposed regions. Some recent image-based NeRF and 3DGS related works utilize multi-exposure LDR images [6, 4, 29] to compensate for the limited dynamic range and model the camera motion [31, 47, 58] to mitigate the influence of motion blur, while they are still susceptible to the requirement of cumbersome multi-exposure capture and severe motion blur under long-exposure conditions.
@@
Event cameras provide a promising sensing modality by asynchronously recording brightness changes in the log domain, thereby preserving high temporal resolution and dynamic range information that is typically lost in conventional frame-based RGB cameras [12]. As shown in the upper half of Fig. 1, event data can effectively record spatial-temporal differential data even in blurred, overexposed, and underexposed areas of an LDR image. Motivated by this advantage, recent approaches have explored using event and RGB image (ERGB) data for the deblurring NVS task [37, 36, 5, 38, 10], while their performance is limited by a lack of grounding in the sensor-physics mismatch between the recorded RGB image and the physical scene brightness. EvHDR-NeRF [6] first incorporates events into the single-exposure HDR and deblurring NeRF reconstruction task. However, its framework largely follows HDR-NeRF [19] by applying the same tone mapping strategy directly to 3D points and extending it to the event-based setting without considering the event generation character.
@@
We argue that achieving accurate and physically consistent 3D representation learning in the ERGB setting requires more than directly extending existing image-based frameworks. In particular, it is essential to explicitly model how scene radiance is transformed into actual sensor measurements through device-specific processes. For RGB cameras, this involves temporal integration and non-linear camera response functions (CRF); for event cameras, this includes contrast thresholding, latency effects, and photometric quantization. Without accounting for these physical transformations, the rendered outputs of the learned model may not align with the supervision signals derived from input images and events, weakening the learning of actual scene geometry and appearance as shown in the lower half of Fig. 1. Therefore, a sensor-physics grounded modeling of both RGB and event sensing pipelines is critical for fully exploiting the complementary information in ERGB data and achieving robust deblurring HDR NVS performance.
@@
In this paper, we present a novel framework that exploits Single-Exposure blurry images and corresponding Events to enable sharp HDR NeRF reconstruction (See-NeRF). Specifically, built upon the physical imaging principles of RGB and event sensors, we adopt NeRF to focus on learning the actual radiance of the scene in the HDR domain and introduce two differentiable CRF models: a 2D pixel-wise RGB CRF model that accounts for exposure integration and camera response function in image generation, and a novel event CRF model that accounts for latency and photometric deviations in event generation. These two modules align the predicted radiance with the sensor measurements and are jointly optimized with the NeRF network under the supervision of input images and events. By modeling the radiance-to-measurement process explicitly, See-NeRF can accurately reconstruct physically consistent, sharp HDR 3D representations, achieving state-of-the-art deblurring HDR NVS results on both public and collected datasets. Our contributions are summarized as follows:
@@
2) We introduce a physically grounded RGB CRF model and a latency-aware photometrically calibrated event CRF model to bridge the discrepancy between physical radiance values and sensor measurements.
3) We collect both synthetic and real datasets for training and testing. The experiment results show that our method achieves the best performance on both ERGB-based HDR and deblurring novel view synthesis tasks.
@@ 2.1 HDR and deblurring novel view synthesis
NeRF [33] extensively promotes the development of the novel view synthesis (NVS) task and 3D implicit representation learning. 3DGS [23] further improve the speed of 3D reconstruction and view rendering. However, NeRF and 3DGS can only learn a sharp LDR scene representation with sharp LDR images. For the HDR NVS task, RawNeRF [32] and RawGS [22] aim to reconstruct an HDR 3D representation from noisy HDR raw images. HDR-NeRF [19] and HDR-GS [4] introduce a 3D points tone mapper to transform the HDR 3D representation into the LDR 3D representation, optimized with multi-exposure LDR images. HDR-HexPlane [50] further extends HDR-NeRF to dynamic scenes. GaussHDR [29] accepts additional context features as input to improve HDR NVS. Gaussian in the dark [53] uses a CNN-based feature map to enhance the tone mapper learning. For the Deblurring NVS task, Deblur-NeRF [31] proposes a deformable sparse kernel module to model the blur kernels. BAD-NeRF [47] and BAD-GS [58] transform the camera poses into SE(3) space and interpolate the poses to recover the camera motion trajectory. DP-NeRF [24] imposes 3D consistency and refines the color with the relationship between depth and blur. Although these image-based methods have achieved promising HDR and deblurring NVS performance, they still fundamentally suffer from the inherent information loss of the motion-blurred LDR images.
@@ 2.2 ERGB-based novel view synthesis
Event cameras are bio-inspired sensors detecting brightness changes at individual pixels with high temporal resolution [12]. This scene radiance measurement paradigm has higher dynamic range and temporal resolution compared to traditional frame-based RGB cameras, attracting a lot of research in the fields of event-based image deblurring [35, 21, 28, 43, 51], HDR image reconstruction [46, 15, 48, 34, 49, 14, 8], and high-speed HDR video generation [39, 61, 52, 59]. Some recent works explore the event-based NVS task [20, 41, 30, 31, 11, 57, 13, 13, 55], while other works integrate both image and event (ERGB) data into the deblurring NVS task, such as NeRF-based works [37, 36, 5, 38, 27, 44] and 3DGS-based works [10, 18, 54, 26]. Addressing the limitations of blurry LDR inputs, EvHDR-NeRF [6] first jointly exploits single-exposure images and events to reconstruct HDR NeRF. However, these previous ERGB-based NVS methods do not fully consider the gap between sensor output and physical radiance, leading to limited HDR and deblurring NVS performance.
@@ 3 Background
In this section, we leverage the principles of event and LDR image generation to explain why events can enhance HDR deblurring NVS with single-exposure blurry LDR images.
@@ 3.1 From scene radiance to LDR images
In an extreme lighting scene with both highlight and low-light areas, a typical RGB camera sensor records the physical scene radiance through a frame-based integral model, generating a raw image with high dynamic range pixel values. To fit human eye perceptual characteristics and reduce storage costs, the raw image is processed by the image signal processor (ISP), applying non-linear ISO gain, RGB demosaicing, tone mapping, gamma compression, white-balance, and quantization to yield an 8-bit LDR image, which is always modeled by a camera response function (CRF) in practice [2]:
where is the raw scene radiance hitting the sensor at time . and are the start and end time of exposure. is the final color value that we can obtain from the LDR image. This integral model and non-linear CRF inherently obscure the temporal dynamics of and compress the critical HDR spatial radiance differential information in highlight and low-light areas, ineluctably leading to artifacts like motion blur, overexposure (pixel values saturation at 255), and underexposure (pixel values clipping at 0) in the LDR images, respectively.
@@ 3.2 From scene dynamics to events
Event cameras asynchronously detect the change of the scene radiance hitting the sensor in the log domain and trigger an event whenever the change reaches the threshold . represents the pixel coordinates and is the trigger time. The change direction is encoded in . This spike model can preserve temporal radiance dynamics and spatial radiance gradient distribution:
which precisely compensates for the temporal and spatial dynamics loss of blurry compressed LDR images (Sec. 3.1).
@@ 3.3 Events for deblurring and HDR NVS
According to Sec. 3.1, the core problem of recovering the sharp HDR scene radiance from blurry LDR images is formulated as the joint estimation of the latent sharp luminance and the CRF . The temporal radiance dynamics of events in Eq. 2 can be used to estimate the latent sharp along the time dimension from blurry as proved in ERGB-based deblurring works [35, 37]. The spatial dynamics of events in Eq. 2 can provide extra radiance differential information for the highlight and low-light areas and expand the dynamic range of the single-exposure LDR image , as shown in Fig. 2, with which the network in Sec. 4.2 can obtain a wider range of as input and infer a CRF curve closer to the ground truth.
@@ 4 Method
In this section, we introduce the See-NeRF and theoretically demonstrate how it leverages sensor-physics priors to achieve high-quality deblurring HDR NVS with single-exposure blurry LDR images and corresponding events
@@ 4.1 HDR scene representation
Firstly, we use the NeRF network with parameters to represent the actual radiance and density of a 3D point in an extreme lighting scene, rather than the camera-processed LDR radiance used in the original NeRF [33]:
@@
Then we use volume rendering to simulate the process of HDR scene rays hitting the imaging sensor at pixel corresponding to the camera pose and obtain the latent sharp HDR radiance at time :
@@
4.2 Physically grounded RGB CRF model
To simulate the integral model of the RGB camera, we discretely sample time points within the exposure time to as in [36], and the predicted raw pixel value of the RGB imaging sensor during exposure is:
where is weight parameters calculated by time points [36]. Since is constantly changing during exposure in the handheld situation, we use the event-guided COLMAP [42] strategy in [37, 36] for the pose estimation. To bridge the gap between the raw sensor value and the LDR image calculated by the CRF in Eq. 1, we used three MLPs to fit this function in the log domain for three color channels separately and obtain the estimated LDR pixels:
@@ 4.2 Physically grounded RGB CRF model
As shown in Fig. 4, unlike HDR-NeRF applying the CRF function directly on the radiance of 3D points in Eq. 3, we employ CRF after volume rendering at the pixel-level for the raw pixel values , aligning with the imaging process in the physical world (Fig. 3), which allows to focus on non-linear color tone mapping learning without being affected by the linear volume rendering, while NeRF focuses on learning 3D density and raw radiance without being affected by . Besides, our blur synthesis on the HDR pixel values (Eq. 6) is more physically grounded. Therefore, our proposed RGB CRF model can significantly improve the HDR and deblurring NVS results as demonstrated by the experiment in Sec. 5.4.
4.3 Latency-aware and photometrically calibrated event CRF model
@@
where means taking the integer between the input and zero that is closest to the input. However, the actual event generation may deviate from this ideal model with a significant delay in low-light regions [17], and due to the fixed contrast threshold in event generation, the minimum detectable brightness change is lower-bounded, leading to quantifiable errors in radiometric estimation. To bridge the gap between actual radiance dynamics and actual event generation, we propose a novel event CRF model comprising three core components: Bayer pattern adaptation, temporal delay compensation, and photometric quantity calibration.
@@
Temporal delay compensation: Since event delay is mainly determined by the pixel radiance values [17], we calculate the delay coefficient with a MLP-based function :
Then, we use a second-order low-pass filter to synthesize the event latency with the estimated coefficient :
We omit in Eq. 10 because the above operation is the same for all pixels. We set and substitute into the event generation model (Eq. 8) to calculate the estimated numbers of events:
Photometric quantity calibration: Since event acquisition exhibits uncertainty when brightness changes are below the threshold or our sampled time points may not precisely align with the timestamp of the last triggered event at this pixel, we pre-estimate an offset from the input events by using , and then subtract this offset for the final estimated number of events to compensate this discrepancy:
@@ 4.4 Loss functions and optimization
We use the LDR image loss and event loss to jointly supervise the optimization of the HDR scene radiance representation (Eq. 3) and the two CRF models:
@@
We take and optimize a coarse and a fine network simultaneously as in [36]. During testing, we can obtain the novel view sharp HDR images with Eq. 6 and LDR images with novel exposure with Eq. 7. We set as a trade-off between performance and training time as in [37]. The implementation details and more evaluation on and are provided in Sec. C and Sec. D of the supplementary.